# ODYSSEE: Oyster Detection Yielded by Sensor Systems on Edge Electronics

### v1

@@ Abstract
Oysters are a keystone species in coastal ecosystems, offering significant economic, environmental, and cultural benefits. However, current monitoring systems are often destructive, typically involving dredging to physically collect and count oysters. A nondestructive alternative is manual identification from video footage collected by divers, which is time-consuming and labor-intensive with expert input.
An alternative to human monitoring is the deployment of a system with trained object detection models that performs real-time, on edge oyster detection in the field. One such platform is the Aqua2 robot. Effective training of these models requires extensive high-quality data, which is difficult to obtain in marine settings. To address these complications, we introduce a novel method that leverages stable diffusion to generate high-quality synthetic data for the marine domain. We exploit diffusion models to create photorealistic marine imagery, using ControlNet inputs to ensure consistency with the segmentation ground-truth mask, the geometry of the scene, and the target domain of real underwater images for oysters. The resulting dataset is used to train a YOLOv10-based vision model, achieving a state-of-the-art 0.657 mAP@50 for oyster detection on the Aqua2 platform. The system we introduce not only improves oyster habitat monitoring, but also paves the way to autonomous surveillance for various tasks in marine contexts, improving aquaculture and conservation efforts.
@@ I INTRODUCTION
Figure 1: Oyster detection system deployed on an Autonomous Underwater Vehicle (AUV). The main image shows oyster detection in real time in a shallow marine environment on the Aqua2 robot. The inset highlights the output bounding boxes around the oysters detected in situ
Oyster reefs are crucial benthic environments, providing numerous ecosystem services, including water filtration, enhanced species richness, shoreline stabilization, and velocity attenuation [1, 2, 3]. Unfortunately, standing oyster stocks in areas such as the Chesapeake Bay [4] and the North Sea [5] have decreased significantly due to over-fishing, global warming, and disease. To mitigate this devastating ecological loss, massive efforts are being made to monitor and restore oyster habitats across Europe [5] and the United States [6, 7, 8, 9].
Despite this effort, the ability to efficiently monitor oyster beds for guided and enhanced restoration remains a major challenge [10, 11, 6]. Traditional monitoring methods, such as those conducted annually in Delaware Bay, New Jersey [12], and biannually in Chesapeake Bay, Maryland [13], rely heavily on destructive sampling techniques (dredging or tonging) and trained technicians to process samples. Among these reports, universal parameters such as “reef areal dimensions, reef height, oyster density, and oyster size-frequency distribution” are commonly assessed [14, 5, 7, 8, 9]. The overall process is costly and labor-intensive, limiting the scalability of surveying efforts.
@@
In fact, by collecting videos or photographs of the beds and assessing the quantity of live and dead oysters, a census of a localized region can be produced to determine the quality of the reef and the harvestable biomass. Similar video-based methods were successfully used to determine habitat provisioning in oyster reefs and aquaculture gear to provide valuable estimates of maxN, species composition, and behavior [19, 20, 21, 22]. However, these methods require highly trained individuals to manually identify and count animals over a large timespan, limiting the depth, scale, and applicability of monitoring for surveying. Autonomous systems can detect and assess live oysters in real time using vision models, offering a promising approach to monitoring. These systems enable on-site decision-making based on live feedback, but training such models requires large datasets of oyster imagery, which are often scarce.
To address this, we leverage diffusion models to generate realistic synthetic images, augmenting real datasets. These combined datasets are used to train models that are then deployed in autonomous underwater vehicles (AUVs) for real-time oyster population surveys. The models trained on these combined datasets are deployed on AUVs, enabling autonomous surveys in oyster fields, and providing accurate counts of oyster populations within their biomes. A key aspect of our approach is that the detection computation is performed in real-time onboard the AUV.
@@ II RELATED WORK
Robotic monitoring of aquatic environments presents technical challenges due to low visibility, complex terrains, and the need for autonomy and real-time decision-making [23, 24, 25]. Edge computing, as a framework that brings data processing applications closer to data sources, can support robot autonomy in environmental monitoring missions[26]. When implemented in autonomous underwater vehicles (AUVs), edge computing allows real-time and local data processing in support of a range of robotic missions, from object detection [27] to autonomous navigation [28] and exploration [29] under resource constraints.
@@
Given the capability of these models to generate representative synthetic data, many have used diffusion models to aid in the training of vision models. He et al. [44] demonstrated the use of synthetic data in training vision models, stating its efficacy in improving the performance of models in zero-shot and low-shot environments. Bansal et al. [45] used synthetic data to achieve higher accuracy with ImageNet models, and Trabucco et al. Off-the-shelf stable diffusion models have also been used [46] to enhance existing data and apply it to few-shot domains. Despite the obvious advantages that diffusion models afford for dataset enhancement, few efforts attempt to the creation of stable diffusion-enhanced datasets for the underwater domain.
Our work combines (i) edge computing for real-time underwater exploration and (ii) advanced synthetic data generation. While previous efforts faced challenges due to small and homogeneous datasets, our approach leverages diffusion models and geometric modeling to create diverse and realistic oyster datasets and deploy state-of-the-art machine learning models such as YOLOv10 [47] running locally on the Aqua2 AUV, ensuring robust performance with edge computation in real-world scenarios. This innovation is crucial to developing robust oyster detection systems and improving the effectiveness of autonomous environmental monitoring in complex marine environments in real-time.
@@ III System Overview
To efficiently monitor oyster habitats in real time, we utilize the Aqua2 robotic platform (Figs 1, 3), designed for nondisruptive environmental monitoring tasks. We will first provide an overview of the Aqua2 robot, followed by a detailed explanation of the software deployed on it. To further address the challenge of limited real-world oyster imagery in reef environments, we supplement our dataset with synthetic data, which will be discussed in later sections of this overview.
@@ III-A Hardware Overview
Aqua2 is a biomimetic hexapedal Autonomous Underwater Vehicle (AUV), designed for minimal environmental disturbance in sensitive habitats such as coral and oyster reefs [48]. Unlike traditional AUVs that use thrust-based propulsion, Aqua2 employs a unique biomimetic reciprocating paddle motion with its six fins, significantly reducing sediment disruption by avoiding thrust wash. This feature is crucial to maintain water clarity in underwater environments, as sediment plumes can obscure sensors and cameras [49]. The Aqua2 vehicle is equipped with a high-performance NVIDIA Jetson Xavier NX processor, which features a 384-core NVIDIA Volta GPU with 48 Tensor Cores, delivering up to 21 TOPS (Tera Operations Per Second) of performance at 15W of power. The processor is complemented by a 6-core NVIDIA Carmel ARM®v8.2 64-bit CPU, providing robust processing capabilities essential for complex, real-time image processing tasks. In addition, the vehicle is equipped with three 2-megapixel cameras optimized for low-light underwater imaging and with scene-based white balancing. The two front E-CAMs capture full HD video at 65 fps, while the rear Blue Robotics camera records at 30 fps. Both have focal lengths of about 3mm, but the front cameras offer a much wider field of view and overall better performance for surveying underwater environments. With electronic on board, Aqua2 can run for five hours on an eight-hour charge of its user-swappable 28.8-V/7.2-Ah lithium-ion battery.
@@ III-B Software Overview
Aqua2’s software architecture leverages the Docker containers for easy management and deployment. NVIDIA’s Container Runtime, compatible with the Open Containers Initiative (OCI), provides GPU access as outlined in III-A. Based on Dustynv/L4t-ml v-r36.2.0111https://hub.docker.com/r/dustynv/l4t-ml container, we integrate Robot Operating System (ROS2), OpenCV, PyTorch v2.2.0 and Ultralytics. The system runs on ROS2 Humble middleware for real-time communication capabilities, modular design, and simplicity. OpenCV handles basic computer vision tasks and Ultralytics installs YOLOv10 and all its dependencies. We developed a custom ROS package that executes the YOLOv10 ROS node to establish the vision pipeline. During operation, Aqua2’s vision pipeline captures images from a downward-facing camera. Pre-trained YOLOv10 submodels, cached on the robot, are quickly loaded to process and compare predictions, highlighting regions of interest and publish this images on the ROS network. These results can be analyzed in real-time by the operator for decision-making or used for autonomous monitoring tasks. With the hardware and software infrastructure in place, we now turn our attention to addressing the challenge of limited real-world oyster imagery by leveraging synthetic data generation to further enhance the model’s performance.
@@ III-C Synthetic Image Generation
Collecting and annotating real images of oyster farms on the seashore is generally tedious, since data collection requires manual robot operation, and annotating often requires expert knowledge. The data collection process is further complicated by limited visibility due to turbulence influenced by weather conditions.
To address these challenges, Stable Diffusion, combined with ControlNet, is leveraged to transfer a synthetic dataset generated in Blender, comprising segmentation masks and depth images, into a dataset that closely resembles a target domain.
This method effectively bridges the gap between synthetic and real-world data, producing images that are more representative of the target domain while maintaining the flexibility and control offered by synthetic data generation. The data generation pipeline is illustrated in Fig. 4.
@@ III-C3 Stable Diffusion and ControlNet
Images generated as detailed in Section III-C1 are provided as the source domain and iteratively selected as ControlNet inputs. Depth maps and segmentation are kept in corresponding pairs such that the segmentation and geometry complement each other. Therefore, guiding the image generation to align with the desired structure and spatial characteristics. Additionally, four randomized reference images from the target domain, collected using the BlueROV robot, are used as image prompts for ControlNet to fine-tune the stylistic and environmental attributes of the generated images. A carefully selected text prompt ensures that Stable Diffusion focuses on the specific target object, namely, the oysters, and it includes negative prompts to mitigate certain biases in the diffusion models. Each generated image from the diffusion model is paired with the corresponding segmentation mask and then used as one of the ControlNet inputs. These masks serve to help generate annotations for the synthetic oyster images, enabling them to act as the ground truth.
@@ IV Experiments And Results
By combining synthetically generated oyster images with real ones, many mixed datasets were prepared with varying proportions of real/synthetic data in the training/validation set. These datasets were then subsequently used to train all of the available YOLOv10 models and comparisons were run between these augmented models and the models which had only seen the real data.
@@ IV-A Training Details
To train the various models, 4000 oyster images were synthetically generated as describe in Section III-C with the size of 640$\times$480, which were then combined with the existing dataset of 2025 real images of 1920$\times$1680 resolution. To this end, we use only 30% of the real dataset for training combined with all the synthetic dataset. We test our method on the remaining 70% of the real dataset to report our result. We train all networks for 300 epochs with an initial learning rate of 0.01 with a cosine annealing learning rate scheduler. The Adam optimizer was used with momentum set to 0.937 and a weight decay of 0.0005. To avoid overfitting. The training batch size was 16, and the model was trained on resized images to 640$\times$640 pixels. In terms of data augmentation, we utilized several techniques, such as flipping, scaling, and random cropping. Close mosaic augmentation was applied for the first 10 epochs, with the mosaic probability set to 1.0. The model was further fine-tuned using automated augmentations, such as randaugment and random erasing with a probability of 0.4. During training, the object detection model was evaluated against a validation set, with an IoU threshold of 0.7. The maximum number of detections per image was capped at 300 to ensure efficient processing of the oyster habitat scenes.
@@ IV-B Experimental Results
Model Inference Time (ms) Frequency (Hz) mAP50 mAP50-95
YOLOv10-N (a) 54.9 9.2 0.458 0.234
YOLOv10-S (b) 71.0 8.9 0.557 0.314
YOLOv10-M (c) 158.1 4.8 0.642 0.468
YOLOv10-B (d) 214.8 3.8 0.645 0.416
YOLOv10-L (e) 264.1 3.2 0.657 0.430
YOLOv10-X (f) 375.2 2.3 0.642 0.424
@@ ACKNOWLEDGMENT
This work was supported by USDA NIFA Sustainable Agriculture System Program under award number 20206801231805, NOAA’s Project ABLE via award number NA22OAR4690620-T1-01, the EU-program EC Horizon 2020 for Research and Innovation under grant agreement No. 101120823, project MANiBOT, and the US Army Corps of Engineers, ERDC Contracting Office under Contract No. W912HZ-22-2-0015. We thank Dr. Christopher Rasmussen and Dr. Arthur Trembanis for their support during the 2024 Autonomous Systems Bootcamp at the University of Delaware, Independent Robotics for assisting with Aqua AUV operations, and Rileigh Hudock, Dr. Edward Hale, and Max Collins for help with data collection.

### v3

@@ Abstract
Oysters are a vital keystone species in coastal ecosystems, providing significant economic, environmental, and cultural benefits. As the importance of oysters grows, so does the relevance of autonomous systems for their detection and monitoring. However, current monitoring strategies often rely on destructive methods. While manual identification of oysters from video footage is non-destructive, it is time-consuming, requires expert input, and is further complicated by the challenges of the underwater environment.
To address these challenges, we propose a novel pipeline using stable diffusion to augment a collected real dataset with photorealistic synthetic data. This method enhances the dataset used to train a YOLOv10-based vision model. The model is then deployed and tested on an edge platform; Aqua2, an Autonomous Underwater Vehicle (AUV), achieving a state-of-the-art 0.657 mAP@50 for oyster detection.
@@ I INTRODUCTION
Figure 1: Oyster detection system deployed on an Autonomous Underwater Vehicle (AUV). The main image shows oyster detection in real-time in a shallow marine environment on the Aqua2 robot. The inset highlights the output bounding boxes around the oysters detected in situ
Oyster reefs are crucial benthic environments, providing numerous ecosystem services, including water filtration, enhanced species richness, shoreline stabilization, and velocity attenuation [1, 2, 3]. Unfortunately, standing oyster stocks in areas such as the Chesapeake Bay [4] and the North Sea [5] have decreased significantly due to over-fishing, global warming, and disease. To mitigate this devastating ecological loss, massive efforts are being made to monitor and restore oyster habitats across Europe [5], Aisa [6, 7] and the United States [8, 9, 10, 11].
Despite this effort, efficiently monitoring oyster beds for guided and enhanced restoration remains a major challenge [12, 13, 8]. Traditional monitoring methods, such as those conducted annually in Delaware Bay, New Jersey [14], and biannually in Chesapeake Bay, Maryland [15], rely heavily on destructive sampling techniques (dredging or tonging) and trained technicians to process samples. Among these reports, universal parameters such as “reef areal dimensions, reef height, oyster density, and oyster size-frequency distribution” are commonly assessed [16, 5, 9, 10, 11]. The overall process is costly and labor-intensive, limiting the scalability of surveying efforts.
@@
In fact, by collecting videos or photographs of the beds and assessing the quantity of live and dead oysters, a census of a localized region can be produced to determine the quality of the reef and the harvestable biomass. Similar video-based methods were successfully used to assess habitat provisioning in oyster reefs and aquaculture gear to provide valuable estimates of relative abundance, species composition, and behavior [21, 22, 23, 24]. However, these methods require highly trained individuals to manually identify and count animals over a large timespan, limiting the depth, scale, and applicability of monitoring for surveying. Autonomous systems can detect and assess live oysters in real time using vision models, offering a promising approach to monitoring. These systems enable on-site decision-making based on live feedback, but training such models requires large datasets of oyster imagery, which are often scarce.
To address this, we leverage diffusion models to generate photorealistic synthetic images, augmenting real datasets. These combined datasets are used to train models that are then deployed in autonomous underwater vehicles (AUVs) for real-time oyster population surveys. The models trained on these combined datasets are deployed on AUVs, enabling autonomous surveys in oyster fields, and providing accurate counts of oyster populations within their biomes. A key aspect of our approach is that the detection computation is performed in real-time onboard the AUV.
@@ II RELATED WORK
Robotic monitoring of aquatic environments presents technical challenges due to low visibility, complex terrains, and the need for autonomy and real-time decision-making [25, 26, 27]. Edge computing, as a framework that brings data processing applications closer to data sources, can support robot autonomy in environmental monitoring missions[28, 29]. When implemented in autonomous underwater vehicles (AUVs), edge computing allows real-time and local data processing in support of a range of robotic missions, from object detection [30, 31] to autonomous navigation [32] and exploration [33] under resource constraints.
@@
Given the capability of these models to generate representative synthetic data, many have used diffusion models to aid in the training of vision models. He et al. [48] demonstrated the use of synthetic data in training vision models, stating its efficacy in improving the performance of models in zero-shot and low-shot environments. Bansal et al. [49] used synthetic data to achieve higher accuracy with ImageNet models, and Trabucco et al. Off-the-shelf stable diffusion models have also been used [50] to enhance existing data and apply it to few-shot domains. Despite the obvious advantages that diffusion models afford for dataset enhancement, few efforts attempt to create stable diffusion-enhanced datasets for the underwater domain.
Our work combines (i) edge computing for real-time underwater exploration and (ii) advanced synthetic data generation. While previous efforts faced challenges due to small and homogeneous datasets, our approach leverages diffusion models and geometric modeling to create diverse and realistic oyster datasets and deploy state-of-the-art machine learning models such as YOLOv10 [51] running locally on the Aqua2 AUV, ensuring robust performance with edge computation in real-world scenarios. The resulting pipeline is crucial to developing robust robust oyster detection systems and improving the effectiveness of autonomous, real-time monitoring of complex marine environments.
@@ III System Overview
To efficiently monitor oyster habitats in real time, we utilize the Aqua2 robotic platform (Figs 1, 3), designed for nondisruptive environmental monitoring tasks. We will first provide an overview of the Aqua2 robot, followed by a detailed explanation of the software deployed on it. To further address the challenge of limited real-world oyster imagery in reef environments, we supplement our dataset with synthetic data, which will be discussed in the section III-C of this overview.
@@ III-A Hardware Overview
Aqua2 is a biomimetic hexapedal Autonomous Underwater Vehicle (AUV), designed for minimal environmental disturbance in sensitive habitats such as coral [52, 53] and oyster reefs [54]. Unlike traditional AUVs that use thrust-based propulsion, Aqua2 employs a unique biomimetic reciprocating paddle motion with its six fins, significantly reducing sediment disruption by avoiding thrust wash. This feature is crucial to maintain water clarity in underwater environments, as sediment plumes can obscure sensors and cameras [55]. The Aqua2 vehicle is equipped with a NVIDIA Jetson Xavier NX processor, which features a 384-core NVIDIA Volta GPU with 48 Tensor Cores, delivering up to 21 TOPS (Tera Operations Per Second) of performance at 15W of power. The processor is complemented by a 6-core NVIDIA Carmel ARM®v8.2 64-bit CPU, providing robust processing capabilities essential for complex, real-time image processing tasks. In addition, the vehicle is equipped with three 2-megapixel cameras optimized for low-light underwater imaging and with scene-based white balancing. The two front E-CAMs capture full HD video at 65 fps, while the rear Blue Robotics camera records at 30 fps. Both have focal lengths of about 3mm, but the front cameras offer a much wider field of view and overall better performance for surveying underwater environments. On a full load, Aqua2 is capable of running for 5 hours in real world environments.
@@ III-B Software Overview
Aqua2’s software architecture leverages the Docker containers for easy management and deployment. NVIDIA’s Container Runtime, compatible with the Open Containers Initiative (OCI), provides GPU access as outlined in III-A. Based on Dustynv/L4t-ml v-r36.2.0111https://hub.docker.com/r/dustynv/l4t-ml container, we integrate Robot Operating System (ROS2), OpenCV, PyTorch v2.2.0 and Ultralytics. The system runs on ROS2 Humble middleware for real-time communication capabilities, modular design, and simplicity. OpenCV handles basic computer vision tasks and Ultralytics installs YOLOv10 and all its dependencies. We developed a custom ROS package that executes the YOLOv10 ROS node to establish the vision pipeline. During operation, Aqua2’s vision pipeline captures images from a downward-facing camera. Pre-trained YOLOv10 submodels, cached on the robot, are quickly loaded to process and compare predictions, highlighting regions of interest and publish these images on the ROS network. These results can be analyzed in real-time by the operator for decision-making or used for autonomous monitoring tasks. With the hardware and software infrastructure in place, we now turn our attention to addressing the challenge of limited real-world oyster imagery by leveraging synthetic data generation to further enhance the model’s performance.
@@ III-C Synthetic Image Generation
Given the challenges of acquiring real-world data, synthetic image generation has become a valuable tool for extending datasets. By using diffusion models with ControlNet, we transform a synthetic dataset, comprising segmentation masks and depth images generated in Blender, into realistic oyster imagery that closely resembles the target underwater environment, enhancing vision model performance. This method effectively bridges the gap between synthetic and real-world data, producing more representative images while retaining the flexibility of synthetic generation. The data generation pipeline in Fig. 4 illustrates the ControlNet inputs on the left side, consisting of target domain images, the synthetic blender images, depth images and segmentation ground truth masks rendered with blender. For further guiadance a positive and negative prompt is provided. The details of the synthetic data generation pipeline are described in the following sections.
@@ III-C3 Stable Diffusion and ControlNet
Images generated as detailed in Section III-C1 are provided as the source domain and iteratively selected as ControlNet inputs. Depth maps and segmentation are kept in corresponding pairs such that the segmentation and geometry complement each other. Therefore, guiding the image generation to align with the desired structure and spatial characteristics. Additionally, four randomized reference images from the real data collection are used as image prompts for ControlNet to fine-tune the stylistic and environmental attributes of the generated images. A carefully selected text prompt ensures that Stable Diffusion focuses on the specific target object, namely, the oysters, and it includes negative prompts to mitigate certain biases in the diffusion models. Each generated image from the diffusion model is paired with the corresponding segmentation mask and then used as one of the ControlNet inputs. These masks serve to help generate annotations for the synthetic oyster images, enabling them to act as the ground truth.
@@ IV Experiments And Results
The mixed datasets obtained via the method described in  III-C were then subsequently used to train all of the available YOLOv10 models and comparisons were run between these augmented models and the models which had only seen the real data.
@@ IV-A Training Details
To train the various models, 4000 oyster images were synthetically generated as described in Section III-C with the size of 640$\times$480, which were then combined with the existing dataset of 2025 real images of 1920$\times$1680 resolution. To this end, we used only 30% of the real dataset for training combined with all the synthetic dataset. We tested our method on the remaining 70% of the real dataset to report our result. We train all networks for 300 epochs with an initial learning rate of 0.01 with a cosine annealing learning rate scheduler. The Adam optimizer was used with momentum set to 0.937 and a weight decay of 0.0005. To avoid overfitting. The training batch size was 16, and the model was trained on resized images to 640$\times$640 pixels. In terms of data augmentation, we utilized several techniques, such as flipping, scaling, and random cropping. Close mosaic augmentation was applied for the first 10 epochs, with the mosaic probability set to 1.0. The model was further fine-tuned using automated augmentations, such as randaugment and random erasing with a probability of 0.4. During training, the object detection model was evaluated against a validation set, with an IoU threshold of 0.7. The maximum number of detections per image was capped at 300 to ensure efficient processing of the oyster habitat scenes.
@@ IV-B Experimental Results
Model Inference Time (ms) Frequency (Hz) $AP_{50}$ mAP50-95
YOLOv10-N 54.9 9.2 0.458 0.234
YOLOv10-S 71.0 8.9 0.557 0.314
YOLOv10-M 158.1 4.8 0.642 0.468
YOLOv10-B 214.8 3.8 0.645 0.416
YOLOv10-L 264.1 3.2 0.657 0.430
YOLOv10-X 375.2 2.3 0.642 0.424
@@ ACKNOWLEDGMENT
This work was supported by USDA NIFA Sustainable Agriculture System Program under award number 20206801231805, NOAA’s Project ABLE via award number NA22OAR4690620-T1-01, the NSF awards 1943205 and 2024741, the EU-program EC Horizon 2020 for Research and Innovation under grant agreement No. 101120823, project MANiBOT, and the US Army Corps of Engineers, ERDC Contracting Office under Contract No. W912HZ-22-2-0015. We thank Dr. Christopher Rasmussen and Dr. Arthur Trembanis for their support during the 2024 Autonomous Systems Bootcamp at the University of Delaware, Independent Robotics for assisting with Aqua AUV operations, and Rileigh Hudock, Dr. Edward Hale, and Max Collins for help with data collection.